OpenStudy (anonymous):

Show that the derivative of f(x) = 3x – 5 is f’(x) = 3. Explain in terms of slope why this is true.

OpenStudy (anonymous):

f'(x) = (3x-5)' = (3x)' - 5' = 3 - 0 = 3. the slope re[resents the geometric interpretation of the derivative.

OpenStudy (anonymous):

okay there is not more information just that?

OpenStudy (anonymous):

the slope is also the tangent of the angle formed by the graph and the ox axis

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i will draw it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The derivative IS the slope, so 3 makes sense is the sloe of the line, if you look at the original

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ignore my typos, you know what i mean lol

OpenStudy (anonymous):

yea i got what you meant

OpenStudy (anonymous):

y intercept (when the grapf intersects the oy axis): x=0 => y=-5 x intercept (when the grapf intersects the ox axis): y=0 => 3x-5=0 => x=5/3 

OpenStudy (anonymous):

so when you want to find the tangent of that angle, which is also the slope you get \[\tan \alpha = \frac{ 5 }{ \frac{ 5 }{ 3 } } = 5 \times \frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

which is 3. your slope.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

you're welcome.